A woman bought some large frames for ​$17 each and some small frames for ​$7 each at a closeout sale. If she bought 23 frames for ​$241​, find how many of each type she bought.
She bought ____ large frames.
She bought ___ small frames

Answer by addingup(3677)

L + S = 23
subtract S from both sides:
L = 23-S
--------------------------
17L + 7S = 241 substitute for L:
17(23-s) + 7S = 241
391 - 17s + 7s = 241
391 - 10S = 241
- 10S = -150 divide both sides by -10, and remember that -/- = + so when you divide just ignore the - signs because your answer will be positive anyway.
s = 15 She bought 15 small frames and
L = 23-S = 23-15 = 8